3. For the network below, find the complex power S for each element. Show that Σs-0. j4 Q 3Ω 50 450 V (ms) -j6 A (ms) HE4. For the network shown, determine (a) the value of the load impedance ZL such that maximum power is delivered to the load a
Answer #1

* Given the corcuit ام v جهان AI Slusov (ms) GLÓA I ams) ✓ Apply ker at node v, V-5445+ I +660°=0. 14 j6 et 5L456 viltu 1/6complexe power supply sy voltage source Svo (5(450) x (16.33|6:22)* -(5445°) x16.33 16.22) = 63.65 +351.14) VA. Sjų = ; x4 x1rotal complex power absorbed = Total complexe power delivered (63.65 2351.14) VA :: ES = 0 → E (Sder - Sass) = 0A Given the circuit بيان 8^ 6604 (rms) jis solo v (rme) —ј6л. find thi Ton To find circuit circuit Zah, voltage source must bfind th:- su vas a o the 6604 & solon jisn -jon Apply Kelat node N Yth - S0 + 4A5 -6.30 (8-j6) jis van leveret } = 6 Vaty [To deliver maximum power to the load, 2 must be equel to 21* 1:22= Zeitalte -(12.41+31-04)* -(12.01 31.06) ::In _Vits Z(+2th